OpenStudy (anonymous):

A car accelerates from 10 km/h to 80km/h in 8.0s. What is the car's acceleration?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Please help!

random231 (random231):

what is the definition of acceleration?

OpenStudy (skullpatrol):

Do you have a definition of acceleration?

OpenStudy (anonymous):

The rate at which velocity changes

random231 (random231):

exactly so what is the change in velocity here in the given time?

OpenStudy (skullpatrol):

$$\huge a=\frac{\delta v}{\delta t}$$

random231 (random231):

so the answer will be (final velocity-initial velocity)/8

OpenStudy (skullpatrol):

$$\Huge \delta v =v_{final}-v_{initial}$$

OpenStudy (skullpatrol):

$$\Huge \delta t = t_{final}-t_{initial}$$

OpenStudy (skullpatrol):

from 10 km/h to 80km/h in 8.0s

random231 (random231):

@OSHelper_Rico arent u getting it?

OpenStudy (skullpatrol):

Watch out for the units of TIME!

OpenStudy (anonymous):

2.4?

OpenStudy (skullpatrol):

What are your units?

OpenStudy (anonymous):

2.4 m/s^2

OpenStudy (anonymous):

10x1000=10000 then 80x1000= 80000. There are 3600 sec in 1 hour. Then 10000/3600= 2.8(simplified), 80000/3600= 22.22.After that 22.22-2.8=19.2/8=2.4.Am I correct or wrong?

random231 (random231):

ure right